# Meta Engine: A Unified Semantic Query Engine on Heterogeneous LLM-Based Query Systems

### v1

@@ 4. Querying Stages and Optimization
In this section we introduce each components and the optimization on them.
@@
4.1. Stage 1: Query Decomposer
You are a database query decomposer.
Your job is to decompose a complex question into
multiple simpler subqueries
that can be answered using the provided multimodal
data sources (table, image, text).
You must follow these formatting rules strictly:
1. The list must contain **at least one** and **at
   most three** subqueries.
2. Each subquery must be a short, clear, and self-
   contained QUESTION.
3. Maintain a logical basic-to-advance reasoning
   order (from basic retrieval to final inference).
We provide 3 types data about several entities
involved in the user query:
1. Table, that contains headers and several rows
2. Image, that either involves the entity or related
   to it.
3. Text, which include detailed description or
   relevent information about the entity.’
Few-Shots Examples
Figure 4. Prompt for Multimodal Query Decomposition
Figure 5. Example of Query Decomposition and operator generation in Meta Engine
4.1.1. Checker and Decomposer
Given a input query, Meta Engine first determines whether it should be answered directly or decomposed into several simpler sub-queries, for which a query complexity checker is invoked. The complexity checker determines whether the question can be answered in a single step or requires multi-step reasoning, with the help of LLM. Specifically, based on an LLM, the checker inspects whether the input query falls in any of the ”Complex Categories” below. If yes, the query is determined to be multi-hop.
  1. (1)
    Multi-part question: multiple tasks, multiple queries, explicit ”and/or/for each”.
  2. (2)
    Multi-step reasoning: needs combining at least 2 distinct facts or sub-results.
  3. (3)
    Aggregation or grouped analysis (sum/avg/count/trend/per-year/per-category).
  4. (4)
    Complex filtering (nontrivial constraints, conditions, temporal scopes).
  5. (5)
    Multi-entity comparison or ranking (compare X vs Y vs Z; find best among many).
  6. (6)
    Pipeline-style tasks (first do X, then do Y using X).
The checker also check whether the query is single-hop by looking for any ”Simple” signs:
  1. (1)
    One clear intent.
  2. (2)
    One entity or one fact.
  3. (3)
    One operator can answer it directly.
  4. (4)
    No decomposition improves clarity.
Particularly, if it is uncertain to the checker, the query is considered as multi-hop by default.
@@ 4.1.1. Checker and Decomposer
If the checker determines the query is single-hop, Meta Engine will bypass the decomposer and treats the original question as a single sub-query. When the query is classified as complex, the query decomposer module is triggered, with the prompt shown in Figure 4. The decomposer reformulates the original question into a set of simpler, self-contained sub-queries that are all single-hop. These sub-queries follow a basic-to-advanced reasoning order, where earlier sub-queries focus on retrieval of basic/grounding facts, and later sub-queries are built upon the results of preceding ones to support higher-level inference. As shown in Figure 5, the original query is decomposed into three sub-queries, where the first retrieves a basic fact about the flag while the last sub-query does advanced, higher-level analytics. Usually each sub-query can be answered within only one modality, e.g., based on one kind of data among text, images or tables.
@@ 4.1.2. Decomposer Optimization
To improve the overall system efficiency, we limit the number of sub-queries decomposed from each query to be at most three. This is because allowing unlimited decomposition may creates a significant number of sub-queries in some datasets. Those sub-queries are often unnecessarily detailed and there are often duplicates or large overlaps between them, which increases latency but does not actually produce better answers. Particularly, many of these sub-queries repeat similar reasoning or retrieve information that is not essential to answering the original query. In addition, long chains of sub-queries make the system more fragile. If an earlier sub-query produces an incorrect or incomplete result, that error will be passed to the later sub-queries and accumulated over time. This leads to slower end-to-end query processing and less accurate final result. By limiting the number of sub-queries, we encourage the system to focus on the most important analysis, reducing unnecessary computation and improving result stability for different datasets.
4.2. Stage 2: Sub-query Routing and Execution
type_of_operators:
imageAnalytics, tableAnalytics, textAnalytics.
You can output multiple possible operators for the
given question.
Each operator must include its type, question, data
sources, and confidence score (between 0 and 1).
Example outputs(multiple operators):
RANKING RUBRIC:
1. Direct Mention or Implication of Modality
- If the question explicitly mentions an image ...
  imageAnalytics highest.
- If the question explicitly mentions a table ...
  tableAnalytics highest.
- If the question asks for textual ... textAnalytics
  highest.
2. Nature of Information Requested
- Quantitative comparison ... tableAnalytics
  priority.
- Historical events ... textAnalytics priority.
- Appearance, colors ... imageAnalytics priority.
3. Ambiguous or Multi-modal Questions
- If multiple operator types could answer, ...
- If ambiguous, ...
4. Confidence Score Ranges (Recommended)
- Highest-ranked operator: 0.70–0.95
- Middle operator: 0.50–0.75
- Lowest operator: 0.30–0.60
The final answer MUST be strictly valid JSON and
follow this schema:
Figure 6. Prompt for Operator Generation and Confidence-Based Ranking
After sub-queries are produced, the next steps are operator generation and routing, which determines how each sub-query will be executed. Figure 5 shows an example on how a user query been decomposed and then further converted to operators. The workflow of operator generation and routing is shown in Figure 3. For each sub-query, the system performs the following steps:
  1. (1)
    Candidate operator generation: Given a sub-query, the operator generator proposes multiple candidate operators, each corresponding to a different modality (TextAnalytics, ImageAnalytics, or TableAnalytics). Instead of using full document contents, the generator relies only on document titles and short captions (e.g., one-sentence summaries for images), to keep routing efficient and low-cost. These candidates are generated by analyzing the intent of the sub-query together with these titles and captions as indicators of the available data sources. Each candidate operator is assigned a confidence score by LLM that estimates how suitable the modality is for answering the sub-query. The prompt is shown in Figure 6.
  2. (2)
    Confidence-based ranking and execution: The candidate operators are sorted in descending order of confidence. The system first executes the highest-ranked operator using its associated data sources. If the execution produces a valid answer, the process terminates and the result is returned. Otherwise, the system proceeds to the next operator in the ranked list. In some cases the top-ranked operator may not be able to properly process the sub-query, so such a progressively attempting is necessary.
  3. (3)
    Iterative fallback and refinement: If no valid answer is obtained by the higher-ranked operator, the system repeats the execution process with lower-ranked operators until an answer is found or all candidates have been exhausted. For sub-queries after the first one, the system additionally refines them before repeating the above steps. Specifically, next sub-query is refined by incorporating execution results from previous sub-queries, such as filling the resolved entity names or intermediate values from the results into next sub-query to reduce ambiguity.
Table 1. Statistics of the datasets
Dataset Queries Text Docs Tables Images Scenario
MultiModalQA 23,817 218,285 10,042 57,058 Multi-hop Q&A
ManyModalQA 1,440 1,440 1,440 1,440 Single-hop Q&A
Text2Vis 1,788 1,788 1,788 1,788 Analytical
FinMMR 1,291 1,291 0 2,269 Analytical
M2QA 1,000 1,884 4,686 37,972 Multi-hop Q&A
@@ 4.3. Stage 3: Aggregator and Final Result
You are an final answer aggregator. Read subqueries
and their tool outputs, then produce the best
possible answer to the original question using ONLY
the provided evidence. Do NOT invent facts.
Goal: Produce an EXTREMELY SHORT final answer based
on all provided Q&As.
Instructions:
1) Analyze all result fields.
2) When multiple answers are present:
    - If one answers the question and the other does
      not...
    - If both answer but differ...
    - If a later sub-query produces an answer that
      logically conflicts with an earlier one...
    - If one answer is descriptive and another...
3) If conflicts remain unresolved, prefer the answer
   that is more specific, complete, and contextually
   appropriate.
STRICT brevity rules for "final_answer":
- HARD CAP: <= 10 words total.
- If yes/no is possible: answer “Yes.” or “No.”
- Always give the most concise factual answer.
Figure 7. Prompt for Final Answer Aggregation with Conflict Resolution
After all sub-queries have been executed, the system put everything into a final answer aggregation module to produce a concise response to the original query from the sub-queries’ results. The inputs to this stage consist of a chain of sub-query–answer pairs, where each sub-query corresponds to an executed operator and its output. The aggregator analyzes all available sub-query results and selects those that directly supports the original query while filtering out irrelevant or non-informative outputs. The prompt is shown in Figure 7. When multiple candidate answers are present, the aggregator resolves conflicts between them by checking logical consistency across the reasoning chain and preferring answers that are coherent with previously established intermediate facts. The final output is aggregated under strict length constraints to encourage direct and unambiguous responses. By explicitly grounding the final answer in intermediate evidence and enforcing consistency across the reasoning chain, this aggregator completes the end-to-end pipeline with a stable and interpretable result.
@@ 4.4. Query Router
In this and the next section we introduce two components, the router and adapters, which are important but not discussed in previous sections in details.
For each operator, we utilized a router to select the most appropriate adapter to execute it. We construct two different types of routers, statistic and learned routers.
@@
Learned router (denoted as MLRouter) is a learning-based, dynamic, sub-query dependent router. It accepts the sub-query, operator, and modality indicators as input, and learns to perform semantic-aware routing based on the sub-query and operator. We implement the learned router using Qwen-0.6B(Team, 2025) as a frozen encoder, followed by a lightweight multilayer perceptron (MLP) classifier. Only the parameters of the MLP are trained, while the backbone language model (Qwen) remains fully frozen throughout training and inference. Router training is conducted offline using historical execution traces collected from all the adapters over the training sub-queries. Specifically, for each sub-query and its operator, all capable adapters are executed individually, and their outputs are evaluated using multiple metrics, including F1, Hit, Coverage, and Semantic Hit (all introduced in Section 5.1). These metric scores are combined into a single quality score by weighted summation, where Semantic Hit and F1 scores are emphasized, while Hit and Coverage serve as secondary criteria. For each training sub-query $q$, the training label is the best adapter $k^{*}_{q}$ for executing this sub-query.
With such training data, the router can learn to choose the proper adapter dynamically based on the sub-query.
The training is performed using a weighted cross-entropy objective, where each training sub-query contributes proportionally to its sample weight. The sample weight for a training sub-query is computed as a function of (1) the absolute quality score of the best-performing adapter and (2) the margin between the best and sub-optimal adapters, as shown in Equation 1 and Equation 2 .
@@ 5.1.2. Methods
DSPY is a programmatic prompting framework that lets users define AI pipelines as modular programs (signatures + modules) and execute them with an underlying LLM/VLM backend. In this baseline,
we instantiate a ChainOfThought module to analyze the original question separately for each modality by conditioning only on the corresponding supporting evidence.
Then, we use a lightweight Predict module as an aggregator to generate the final response.
LlamaIndex is a programming framework for building LLM-based workflows over heterogeneous data sources. In this baseline, we use its indexing and querying abstractions to implement a retrieval-augmented generation (RAG) workflow for text and table modalities, and a direct vision-language querying workflow for images.
The final answer is generated by synthesizing the modality-level outputs with an LLM.
Meta-Engine, as we mentioned in the previous parts,
we propose two variants under the same framework: Meta-Engine (MLRouter) and Meta-Engine (StatRouter). Both use identical components and configurations for complexity checking, decomposition, execution and aggregation, differing only in the routing strategy. MLRouter is the learned router, while StatRouter is the statistic router.
@@ 5.1.3. Metrics
We evaluate end-to-end answer quality using both token-based metrics and semantic similarity metrics. For datasets that provide multiple acceptable ground-truth answers per question, we compute each metric between the prediction and every ground-truth answer, and report the maximum score to account for answer variability.
Let $\hat{y}$ denote the predicted answer and $y$ is a ground-truth answer. After normalization (lowercasing, punctuation removal, and whitespace trimming), we tokenize both answers into token multisets (i.e., bags of tokens preserving multiplicity). Let T(y^)T($\hat${y}) and $T(y)$ denote the resulting multisets, and let
(3) $m(\hat{y},y)=|T(\hat{y})\cap T(y)|$
be the multiset overlap count.
Token-level F1.
We compute precision and recall from token overlap:
(4) $\mathrm{Prec}(\hat{y},y)=\frac{m(\hat{y},y)}{|T(\hat{y})|},\qquad\mathrm{Rec}(\hat{y},y)=\frac{m(\hat{y},y)}{|T(y)|},$
and the token-level F1 score:
(5) $\mathrm{F1}(\hat{y},y)=\frac{2\cdot\mathrm{Prec}\cdot\mathrm{Rec}}{\mathrm{Prec}+\mathrm{Rec}}.$
Hit and Semantic Hit.
To complement strict lexical matching, we report two overlap-based metrics. Hit is a binary metric indicating whether $\hat{y}$ shares at least one token with $y$:
(6) $\mathrm{Hit}(\hat{y},y)=\mathbb{I}[m(\hat{y},y)>0].$
However, token-overlap metrics may penalize valid paraphrases or fail to capture subtle factual errors. To obtain a higher-level semantic assessment, we additionally employ an LLM as a semantic hit judge. Given the question, the prediction $\hat{y}$, and the set of ground-truth answers to the question, the judge determines whether $\hat{y}$ conveys the same factual content with at least one ground-truth answer, ignoring surface-form variations such as paraphrasing or reordering. If yes, $\hat{y}$ is a semantically correct answer, and its Semantic Hit is 1, as shown below:
(7) $\mathrm{Sem\_Hit}(\hat{y},y)=\mathbb{I}[\hat{y}\ is\ semantically\ equivalent\ to\ y].$
In the evaluation results, we report the average Hit and average Semantic Hit, i.e., the fraction of hit questions over all questions, and the fraction of semantically hit questions over all.
In addition, we also measure the end-to-end query processing time as efficiency metric.

### v2

@@ 4. Querying Stages and Optimization
In this section we introduce details of each components and the optimization on them.
@@
4.1. Stage 1: Query Preprocessing
Algorithm 1 Dynamic Query Decomposition
1:Query $q$, flag for complexity checking $F_{c}$, maximum subuqery number max_outmax\_out, maximum iterations max_itmax\_it
2:Set of subqueries $\mathcal{S}$, Dependency graph $\mathcal{G}_{S}$
3:iter := 0
[…]
5:while $iter\leq max\_it$ do
6:  if $F_{c}=True$ then
7:    if Check($q$) = ”Complex” then
8:     $\mathcal{S}:=\texttt{Decompose}(q)$
9:    else
10:     $\mathcal{S}:=\{q\}$
11:  else
12:    $\mathcal{S}:=\texttt{Decompose}(q)$
13:  if $|\mathcal{S}|>max\_out$ then
14:    iter := iter + 1
15:    Set feedback, emphasizing subquery number
16:  else
17:    $\mathcal{G}_{S}$ := ParseDependency($\mathcal{S}$)
18:    return $\mathcal{S}$, $\mathcal{G}_{S}$
[…]
20:return $\mathcal{S}$, $\mathcal{G}_{S}$
4.1.1. Query Decomposer
Complex queries are hard to processed accurately, therefore Meta Engine provides a dynamic query decomposition method to split the input query into several simple subqueries. The method involves two modules, query complexity checker and the decomposer. As described in Algorithm 1, given an input query, the complexity checker first determines whether the query is complex enough, and for a complex query (e.g., multi-hop queries), the decomposer will split it into a set of simpler sub-queries, each corresponding to a specific task within a single data modality. If the checker determines the query is already simple (like single-hop), Meta Engine will bypass the decomposer and treats the original query as a single sub-query. Users can turn off the checking, then all queries will be decomposed.
Both checker and decomposer are built on top of LLM. The checker inspects whether the input query falls in any of the ”Complex Categories”, like multi-step reasoning, aggregation or groupby analytics, complex filtering, etc. It also looks for any ”Simple” signs in the query, like whether the query includes one clear intent or asks for only one entity or fact. Particularly, if it is uncertain to the checker, the query is considered as complex by default.
@@ 4.1.1. Query Decomposer
The resulted sub-queries follow a basic-to-advanced reasoning order, where earlier sub-queries focus on retrieval of basic/grounding facts, and later sub-queries are built upon the results of preceding ones to support higher-level inference. For example, in Figure 5, the original query is decomposed into three sub-queries, where the first retrieves a basic fact about the flag while the last sub-query does advanced, higher-level analytics. In addition, the decomposer will ensure each sub-query can be answered within only one modality, i.e., based on one kind of data among text, images or tables. This is critical to guarantee that each subquery can be processed by one data operator.
@@ 4.1.2. Decomposer Optimization
To improve the overall system efficiency, we design two optimizations: (1) limiting the number of sub-queries generated, and (2) dependency-based subquery parallelism.
@@
Allowing unlimited decomposition may creates a significant number of sub-queries in some datasets. Those sub-queries are often unnecessarily detailed and there are often duplicates or large overlaps between them, which increases latency but does not actually produce better answers. So we allow users to specify the maximum number of subqueries they want, by the parameter max_outmax\_out in Algorithm 1. The query decomposition will be redo if the generated subqueries exceed max_outmax\_out. In our evaluation, this redo rarely happens, therefore it does not slow down the decomposition. By limiting the number of sub-queries, we encourage the pipeline to focus on the most important analysis, reducing unnecessary computation and improving result stability for different datasets.
Another optimization is a dependency graph of subqueries, which will be used later to parallelize the subquery execution. Given the basic-to-advanced decomposition, later sub-queries may depend on the results of earlier ones. But the subqueries are usually not fully dependent. For example, a query asking “Which person has a lower salary in 2025: the person in blue with a dog in the picture XXX, or the person with the highest-level position in company YYY” could be decomposed into three subqueries: (1) ”What is the salary of the person in blue with a dog in the picture XXX”, (2) ”What is the salary of the person with the highest-level position in company YYY”, and (3) ”Which person has a lower salary between them”. In this example, the subquery 3 relies on subqueries 1 and 2, while subqueries 1 and 2 have no dependency between each other. Such partially independent subqueries can be parallely executed to improve the pipeline efficiency. To handle such dependencies, the decomposer further organizes the subqueries into a dependency graph whose nodes are subqueries and edges are dependency between subqueries, which is used in execution stage.
Original Query: “When was the last time capital punishment took place in the state whose flag features a bear and has a major city in the Sun Belt?”  Sub-query 1 [image]
“Which state has a flag that features a bear?”
$\rightarrow$
imageAnalytics("Which state has a flag that features a bear?", [’California’])
Result: California
Sub-query 2 [table] $\leftarrow$ depends on Sub-query 1
“Does this state have a major city located in the Sun Belt region?”
$\rightarrow$
tableAnalytics("Does California have a major city located in the Sun Belt region?", [’Sun Belt’])
Result: Anaheim, Bakersfield, Fresno, Long Beach, Los Angeles, Oakland, Riverside, Sacramento, San Bernardino, San Diego, San Jose, San Francisco
Sub-query 3 [text] $\leftarrow$ depends on Sub-query 1 and 2
“When was the last time capital punishment took place in this state?”
$\rightarrow$
textAnalytics("When was the last time capital punishment took place in California?", [’Capital punishment in California’])
Result: 2006 $\Rightarrow$ Final Answer: 2006 in California.
Figure 5. An example of subquery decomposition and dependency parsing, operator selection, execution.
4.2. Stage 2 & 3: Query Planning and Execution
After sub-queries are produced, the next stages are logical planning, physical planning and execution, which is shown in Figure 4.
In Meta Engine, the logical planning is conducted by a confidence-based operator ranking. As mentioned in Section 3.1.2, it is hard to accurately predict which data modality includes the necessary information to process a subquery. So instead of predicting the most proper data operator, Meta Engine ranks the available data operators (TextAnalytics, ImageAnalytics, and TableAnalytics in this paper) by prediction confidence, and execute them progressively later. Instead of using full data contents, this ranking relies only on metadata like document titles and short captions (e.g., one-sentence summaries for images), to keep the planning efficient and low-cost. Given the metadata and subquery, each candidate operator is assigned a confidence score (from 0 to 1) by the LLM-based Operator Generator, which estimates how suitable the modality and operator is for processing the subquery. The ranking will consider whether the subquery has direct mention or implication of the modality, whether the requested information refers to any specific modality, and so on. The scored candidate operators are sorted in descending order of confidence.
Then in physical planning, each data operator will be mapped to a specific adapter’s interface by the query router. The query router intelligently maps each operator to the most appropriate underlying specialist system. Note that each operator could have multiple choices, like LOTUS, DSPy and VLM can all execute ImageAnalytics. In these cases the router will predict the best-performing adapter on the given operator and subquery. For this goal, we design two types of routers: a learned router which is a query-dependent classifier on top of a pre-trained LLM, and a statistic router that is query-agnostic, purely based on history statistics. Our evaluation shows the learned router is overall more accurate, but statistic router outperforms it in some cases. We provide more details in Section 4.4.
After the physical planning, each candidate operator in the ranked list has an adapter to execute it. There are two levels of execution, operator execution and subquery execution. At operator level, the ranked operators will be executed by their adapters progressively: they are executed in order (higher confidence operator first). If an operator does not return valid/meaningful result, next operator will be executed, until meaningful result is returned. This progressive execution ensures the processing quality, since the top-ranked operator may not always be the most proper one for the subquery. Figure 5 shows an example on the operator that returns valid results for each subquery. In our evaluation, the average number of executed operators per subquery ranges from 1.05 to 1.2, proving that this progressive execution raises only negligible overhead.
Executing a subquery is essentially executing its data operators as above. Additionally, at subquery level, the execution is optimized with two techniques: iterative subquery refinement and dependency-based parallelization. The dependency graph generated in decomposition is used by both techniques. First, for sub-queries after the first one, Meta Engine additionally refines them before execution. Specifically, next sub-query is refined by incorporating execution results from previous sub-queries, such as filling the resolved entity names or intermediate values from the results into next sub-query to resolve abstract references into concrete entities and reduce ambiguity. Before executing the $i$-th sub-query, the results from sub-queries $1$ through i1i{-}1 which the $i$-th sub-query has dependency to are provided as additional context. With such additional context, the $i$-th sub-query will be refined as above by LLM. Figure 5 shows an example on how the subqueries are refined, e.g., the result “California” from subquery 1 is filled into subquery 2, shown in its TableAnalytics operator. By such a dynamic refinement, the sub-queries are stated more clearly and the system avoids wasting time in duplicate computation across sub-queries
Second, execution of subqueries is parallelized based on their dependencies. Given the dependency graph, Meta Engine first conduct a BFS-like search to split the graph into stages based on the depth, where each subquery in a stage is dependent to some subqueries in previous stages, and all subqueries in the same stage are independent from each other. Based on such split, the subqueries are executed sequentially across stages while parallelly within each stage. Such a dependency-based parallelization ensures a competitive execution efficiency given the relatively slow LLM calls.
@@ 4.3. Stage 3: Aggregator and Final Result
This final component aggregates the individual pieces of information and generates the final results. It refines the results from different subqueries (like deduplicating the overlaps between results) and synthesize them to the comprehensive and accurate final result. Ideally, the result of the last subquery is the final result to the original query. But there could be information conflicts between the results of last subquery and previous subqueries. So a major responsibility of the aggregator is to resolve conflicts between them by checking logical consistency across the reasoning chain and preferring the result that are coherent with previously established intermediate facts. This aggregator is optional in a pipeline. Without it, the final result will be the result from the last subquery by default.
@@ 4.4. Query Router
In this and the next section we introduce two components, the router and adapters. For each operator, we utilized a router to select the most appropriate adapter to execute it. We construct two different types of routers, statistic and learned routers.
@@
Learned router (denoted as MLRouter) is a learning-based, dynamic, sub-query dependent router. It accepts the sub-query, operator, and modality indicators as input, and learns to perform semantic-aware routing based on the sub-query and operator. We implement the learned router using Qwen-0.6B(Team, 2025) as a frozen encoder, followed by a lightweight multilayer perceptron (MLP) classifier. Only the parameters of the MLP are trained, while the backbone language model (Qwen) remains fully frozen throughout training and inference. Router training is conducted offline using historical execution traces collected from all the adapters over the training sub-queries. Specifically, for each sub-query and its operator, all capable adapters are executed individually, and their outputs are evaluated using multiple metrics, including F1, Hit, Coverage, and Semantic Hit (all introduced in Section 5.1). These metric scores are combined into a single quality score by weighted summation, where Semantic Hit and F1 scores are emphasized, while Hit and Coverage serve as secondary criteria. For each training sub-query $q$, the training label is the best adapter $k^{*}_{q}$ for executing this sub-query. With such training data, the router can learn to choose the proper adapter dynamically based on the sub-query. The training is performed using a weighted cross-entropy objective, where each training sub-query contributes proportionally to its sample weight. The sample weight for a training sub-query is computed as a function of (1) the absolute quality score of the best-performing adapter and (2) the margin between the best and sub-optimal adapters, as shown in Equation 1 and Equation 2 .
@@ 4.5. Adapters
Table 2. Statistics of the datasets
Dataset Queries Text Docs Tables Images Scenario
MultiModalQA 23,817 218,285 10,042 57,058 Multi-hop Q&A
ManyModalQA 1,440 1,440 1,440 1,440 Single-hop Q&A
Text2Vis 1,788 1,788 1,788 1,788 Analytical
FinMMR 1,291 1,291 0 2,269 Analytical
M2QA 1,000 1,884 4,686 37,972 Multi-hop Q&A
@@ 5.1.2. Methods
DSPY is a programmatic prompting framework that lets users define AI pipelines as modular programs (signatures + modules) and execute them with an underlying LLM/VLM backend. In this baseline, we instantiate a ChainOfThought module to analyze the original question separately for each modality by conditioning only on the corresponding supporting evidence. Then, we use a lightweight Predict module as an aggregator to generate the final response.
LlamaIndex is a programming framework for building LLM-based workflows over heterogeneous data sources. In this baseline, we use its indexing and querying abstractions to implement a retrieval-augmented generation (RAG) workflow for text and table modalities, and a direct vision-language querying workflow for images. The final answer is generated by synthesizing the modality-level outputs with an LLM.
Meta-Engine, as we mentioned in the previous parts, we build two pipelines with Meta Engine framework: Meta-Engine (MLRouter) and Meta-Engine (StatRouter). Both use all the components and identical configurations (as the Pipeline 1 in Figure 3), differing only in the routing strategy. MLRouter is the learned router, while StatRouter is the statistic router.
@@ 5.1.3. Metrics
We evaluate end-to-end answer quality using token-based and semantic similarity metrics. For datasets with multiple acceptable ground-truth answers per question, we compute each metric against every ground-truth answer and report the maximum score.
Let $\hat{y}$ denote the predicted answer and $y$ a ground-truth answer. After normalization (lowercasing, punctuation removal, whitespace trimming), we tokenize both into token multisets T(y^)T($\hat${y}) and $T(y)$, and write $m(\hat{y},y)=|T(\hat{y})\cap T(y)|$ for the multiset overlap count. We compute the standard token-level F1 score from the token-overlap precision and recall. We also use Hit and Semantic Hit. Hit is a binary indicator of whether $\hat{y}$ shares at least one token with $y$, i.e., $\mathrm{Hit}(\hat{y},y)=\mathbb{I}[m(\hat{y},y)>0]$. However, token-overlap metrics may penalize valid paraphrases or miss subtle factual errors. We therefore additionally employ an LLM as a semantic judge: given the question, $\hat{y}$, and the ground-truth answer set, the judge determines whether $\hat{y}$ conveys the same factual content as at least one ground-truth answer, ignoring surface-form variations. We define $\mathrm{Sem\_Hit}=1$ if the judge deems $\hat{y}$ semantically equivalent, and 0 otherwise. We report average Hit and average Semantic Hit (i.e., fraction of questions achieving a hit) across all questions. We also measure end-to-end query processing time as an efficiency metric.